# Continuous K-space Recovery Network with Image Guidance for Fast MRI Reconstruction

### v1

@@ III-B5 Loss Functions and Training Strategy
Following [60, 42], we evaluate the intermediate outputs of various modules by using the $L_{2}$ loss, which consists of two terms. To detail, the first term is used to constrain the consistency between the reconstructed k-space $\widehat{\mathrm{K}}_{i}$ and corresponding fully sampled k-space, i.e., $K_{lr}$ or $K$, while the second term is used to constrain the consistency between the reconstructed image $\widehat{\mathrm{I}}_{i}$ and the corresponding fully sampled image, i.e., $I_{lr}$ or $I$:
@@
where $I$, $I_{lr}$, and $K_{lr}$ is obtained via $I=IFFT(K)$, $I_{lr}=downsample(I)$, and $k_{lr}=FFT(I_{lr})$, respsectively. Furthermore, to mitigate the issues of over-smoothing and distortion and obtain finer reconstruction results step-by-step, we design a multi-stage training strategy, as illustrated by Algorithm 1. Specifically, we divide the training process into four stages, i.e., $stage=stage\_i,when{{\color[rgb]{1,1,1}1}}epoch\in\left[E_{i-1},E_{i}\right]|_{i=1}^{i=4}$. In summary, our IGIT-Net recovers the precise MRI images progressively by optimizing the loss at various stages:
$L_{stage\_j}={\textstyle\sum_{i=1}^{i=j}L_{i}}$ (18)
IV EXPERIMENTS
IV-A Implementation Details
IV-A1 Datasets
We use two real k-space datasets, CC359 and fastMRI, and one simulated dataset, IXI, as described below.
@@ IV-A1 Datasets
CC359 dataset The public brain MR raw data set—the Calgary-Campinas dataset, which comes from a clinical MR scanner (Discovery MR750; GE Healthcare, Waukesha, WI, USA)—is used to train and test our proposed model. In total, $4129$ slices from $25$ volumes are randomly selected to form the training set, and the testing set is composed of $1650$ slices from $10$ other volumes.
@@
fastMRI dataset The fastMRI dataset is currently the largest open-source MRI dataset with raw fully sampled k-space data, which includes a single-coil knee dataset and a multi-coil brain dataset. The single-coil knee dataset which consists of $973$ training volumes ($98772$D slices) and $199$ validation volumes ($6140$ 2D slices) is employed in the experiments to demonstrate the effectiveness of our IGKR-Net.
@@
IXI dataset The IXI dataset consists of $574$ brain MRI volumes from normal and healthy subjects collected from three different hospitals in London. After pre-processing, $46226$ 2D slices from $368$ volumes are used for training, $11562$ 2D slices from $92$ volumes are used for validation, and $14315$ 2D slices from $114$ volumes are used for testing. Since the IXI dataset does not contain the original k-space data, we use the Fast Fourier Transform (FFT) to obtain the simulated k-space data.
IV-A2 k-space undersampling
@@
For these datasets, two different types of masks are tested, i.e., 1D Cartesian undersampling mask and 2D Gaussian undersampling mask. For each sampling, $20\%$ and $40\%$ undersampling rates are tested. Following [26], all undersampling masks are generated by first including some number of adjacent lowest frequency k-space lines or points to provide a fully-sampled k-space region. Specifically, for the 1D Cartesian undersampling mask, the fully-sampled central region includes $8\%$ of all k-space lines; for the 2D Gaussian undersampling mask, $16\%$ of all k-space points are included. Then, the remaining k-space lines (for 1D Cartesian mask) or points (for 2D Gaussian mask) are included uniformly at random, with the probability set so that, on average, the undersampling mask achieves the desired acceleration factor.
IV-A3 Baselines and Training Details
@@
We implement our model in PyTorch using Adam with an initial learning rate of $31e-$ 3 and train it on $1$ NVIDIA A100 GPU with $40$GB of memory. The number of epochs $(E_{0},E_{1},E_{2},E_{3},E_{4})$ for multi-stage training are set as $(0,20,60,100,200)$.
We compare our model with the following algorithms. Zero-Filling (ZF); ResNet[56], which is trained using Adam with a learning rate of $31e\-/$ 3; the end-to-end architecture, UNet[26], provided by fastMRI, which is retrained using Adam with a learning rate of $31e\-/$ 3; RefineGAN[57], a deep learning-based generative adversarial model, which is retrained using its default parameter settings; SwinMR[36], a swin transformer based method, which is trained for $100,000$ steps using Adam optimizer and the initial learning rate is set to $42e\-/$ 4 and decayed by $0.5$ every $10,000$ steps from the $50,000$th step; SwinGAN[58], which combines the swin transformer and generative adversarial network and is retrained using its default settings; OCUCFormer[59], an over-complete under-complete transformer network, which is trained using the Adam optimizer with an initial learning rate of $31e\-/$ 3.
@@
For the quantitative study, Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Normalized Mean Squared Error (NMSE), and Learned Perceptual Image Patch Similarity (LPIPS) are utilized for evaluation.
IV-B Experimental Results
IV-B1 Comparisons with other methods
In this section, we compare the performance of our IGKR-Net with other methods under different sampling patterns and acceleration rates on three public datasets.
@@ IV-B1 Comparisons with other methods
The quantitative results of comparisons on different datasets are shown in Table I, Table II, and Table III, respectively. The best result in each column is highlighted in bold. Based on these results, we can find that: (1) Accelerated sampling leads to zero-filled MRI images with extremely poor quality; (2) Classical deep learning methods, such as UNet and ResNet, reduce image artifacts to some extent, but the restoration results remain suboptimal; (3) Current state-of-the-art methods, such as RefineGAN, SwinMR, and OCUCFormer, have significantly improved image reconstruction quality, but their performance degrades on challenging tasks with high acceleration rates (sampling rate = 20%). (4) Our method consistently achieves the best performance with significant improvements on all datasets under different sampling scenarios. For example, as shown in Table I, our method shows the superiority of $1.54$ dB and $2.63$ dB in PSNR over OCUCFormer under 20% and 40% 2D random sampling masks on the CC359 dataset, respectively. As shown in Table III, our model outperforms OCUCFormer by $0.43$ dB and $0.22$ dB in PSNR under 20% and 40% 1D random masks on the IXI dataset, respectively.
The qualitative comparison results are shown in Fig.2, including the reconstructed MRI images and absolute differences (10×10$\times$10 ×) of standardized pixel intensities between reconstructed results and Ground Truth (GT) of different methods under various masks. As observed, reconstruction using zero-filling (ZF) at high acceleration rates leads to significant aliasing artifacts and loss of anatomical details. In comparison to ZF, deep learning methods based on CNNs such as ResNet and UNet show improvement in reconstruction but still suffer from severe edge blurring and substantial detail loss. SwinGAN and OCUCFormer partially alleviate these issues, yet in challenging tasks with high acceleration rates, these methods lose some crucial anatomical details. In contrast, our IGKR-Net demonstrates robustness to sampling patterns and acceleration rates. Leveraging information from the image domain to guide the recovery in k-space, our method better preserves important anatomical details, as indicated by the arrows and ellipses.
@@
Figure 5: Results of the experiments under masks on CC359 dataset. The X-axis represents the sampling ratio, while the Y-axis corresponds to various metrics.
TABLE IV: Quantitative results of the experiments on different masks.
@@
ratio Method PSNR↑ SSIM↑ NMSE↓ LPIPS↓ PSNR↑ SSIM↑ NMSE↓ LPIPS↓
Unet 25.72 0.8010 0.0352 0.1082 26.50 0.7909 0.0301 0.1075
OCUCFormer 20.80 0.5677 0.3890 0.2839 26.27 0.7363 0.0327 0.0795
10% Ours 27.11 0.8262 0.0267 0.0960 28.96 0.8424 0.0181 0.0719
Unet 29.34 0.8756 0.0156 0.0618 28.34 0.8417 0.0197 0.0769
OCUCFormer 29.96 0.8888 0.0303 0.0395 30.10 0.8554 0.0324 0.0399
20% Ours 33.06 0.9234 0.0069 0.0362 34.26 0.9258 0.0058 0.0298
UNet 30.15 0.8940 0.0127 0.0532 29.73 0.8738 0.0147 0.0564
OCUCFormer 33.84 0.9286 0.0152 0.0174 33.62 0.9110 0.0162 0.0226
30% Ours 35.35 0.9530 0.0039 0.0213 38.38 0.9600 0.0024 0.0147
UNet 31.73 0.9189 0.0090 0.0408 30.59 0.8935 0.0122 0.0481
OCUCFormer 38.57 0.9656 0.0058 0.0077 36.58 0.9396 0.0092 0.0131
40% Ours 38.78 0.9703 0.0020 0.0143 41.22 0.9753 0.0013 0.0085
Unet 34.08 0.9442 0.0053 0.0276 32.06 0.9160 0.0087 0.0402
OCUCFormer 39.68 0.9746 0.0047 0.0063 38.84 0.9571 0.0056 0.0086
50% Ours 43.34 0.9844 0.0008 0.0060 46.04 0.9875 0.0005 0.0042
TABLE V: Quantitative results of the ablation studies and analysis using different masks on the CC359 validation dataset.
1D Cartesian, ratio = 20% 1D Cartesian, ratio = 40%
Method PSNR↑ SSIM↑ NMSE↓ LPIPS↓ PSNR↑ SSIM↑ NMSE↓ LPIPS↓
W/O DIFM 30.42 0.8826 0.0120 0.0612 33.67 0.9342 0.0056 0.0373
W/O TARM 31.97 0.9091 0.0086 0.0440 37.25 0.9634 0.0026 0.0186
ours 33.06 0.9234 0.0069 0.0362 38.78 0.9703 0.0020 0.0143
2D Gaussian, ratio = 20% 2D Gaussian, ratio = 40%
W/O DIFM 32.31 0.8838 0.0079 0.0733 37.34 0.9494 0.0025 0.0291
W/O TARM 36.11 0.9394 0.0036 0.0262 42.36 0.9788 0.0009 0.0058
Ours 37.09 0.9512 0.0029 0.0184 43.04 0.9807 0.0008 0.0048
IV-B2 Experiments on masks
This experimental study aims to evaluate the performance of IGKR-Net using different undersampling trajectories. Five 1D Cartesian undersampling trajectories including 1D Cartesian $10\%$, 1D Cartesian $20\%$, 1D Cartesian $30\%$, 1D Cartesian $40\%$ , and 1D Cartesian $50\%$, as well as five 2D Random undersampling trajectories including 2D Random $10\%$, 2D Random $20\%$, 2D Random $30\%$, 2D Random $40\%$, and 2D Random $50\%$ are applied in this experiment. This experiment compares the SSIM, PSNR, NMSE, and LPIPS of UNet, OCUCFormer, and our IGKR-Net, and is conducted using the CC359 dataset.
The quantitative results of the experiment on different masks are shown in Table IV and Fig. 5, respectively. According to the results, we find that: (1) As shown in Table IV, when using the same type of mask, e.g., 1D Cartesian mask, our IGKR-Net consistently demonstrates significant improvements across different sampling rates compared to other methods. For example, our model outperforms refineGAN by $3.66$ dB and $0.21$ dB under 1D Cartesian mask at $50\%$ and $40\%$ sampling rates, respectively. (2) As shown in Fig.5, our method’s performance under 2D masks (denoted as solid lines) even consistently surpasses that of other methods under 1D masks (denoted as dashed lines ) in various metrics across different sampling rates, despiting 1D mask being recognized as a simpler task. (3) OCUCFormer achieves satisfactory results at higher sampling rates, while its performance significantly declines under more challenging lower rates. This indicates that this method is less suitable for challenging tasks while our method shows robustness across various masks.
@@ IV-B2 Experiments on masks
The sample of reconstructed images and absolute differences of standardized pixel intensities (10×10$\times$10 ×) between reconstructed images and GT are shown in Fig.6. According to the results, it can be observed that as the k-space sampling ratio increases, the performance of different methods steadily improves. It is noteworthy that in challenging tasks with low sampling rates, e.g., 1D or 2D $10\%$, other methods obtain extremely poor performance, while our method can still reconstruct relatively clear contour edges. Overall, our IGKR-Net achieved a higher reconstruction quality compared to other methods under various undersampling trajectories.
Figure 6: Visual comparison of the experiment under different masks.
IV-C Ablation Studies and Analysis
IV-C1 Efficacy of Key Components
To validate the effectiveness of our proposed Image Domain Guidance Module (IDGM) and Tri-Attention Refinement Module (TARM), ablation experiments are conducted on these two modules. Specifically, we design two network variants: (1) w/o DIFM, in which the DIFM module is removed from the network and the k-space spectrum outputted from the LRIT is directly used for HRIT; (2) w/o TARM, the TARM module is removed from the network and the MRI image outputted from the HRIT is regarded as final reconstruction result.
Figure 7: Visual comparison with k-space error maps under different masks.
The performance of the above variant networks is tested by using different masks on the CC359 dataset, and the experimental results are shown in Table V. It can be observed that after removing the TARM module, the performance decreased significantly, which verifies that our proposed TARM can better refine the output and thus guide MRI image reconstruction. In addition, the W/O IDGM network variant achieves the poorest results, which demonstrates that our proposed IDGM module can mine semantic information from undersampled MRI images to obtain more informative outputs.
IV-C2 Efficacy of recovering k-space

### v2

@@ 3.2.5 Loss Functions and Training Strategy
Following Zhou and Zhou (2020); Zhao et al. (2023), we evaluate the intermediate outputs of various modules by using the $L_{2}$ loss, which consists of two terms. To detail, the first term is used to constrain the consistency between the reconstructed k-space $\widehat{{K}}_{i\in\left\{1,2,3,4\right\}}$ and corresponding fully sampled k-space, i.e., $K_{lr}$ for $\widehat{{K}}_{1}$ and $K$ for $\widehat{{K}}_{i\in\left\{2,3,4\right\}}$. The second term is used to constrain the consistency between the reconstructed image $\widehat{{I}}_{i\in\left\{1,2,3,4\right\}}$ and the corresponding fully sampled image, i.e., $I_{lr}$ for $\widehat{{I}}_{1}$ and $I$ for $\widehat{{I}}_{i\in\left\{2,3,4\right\}}$:
@@
where $I$, $I_{lr}$, and $K_{lr}$ is obtained via $I=IFFT(K)$, $I_{lr}=downsample(I)$, and $k_{lr}=FFT(I_{lr})$, respsectively. Furthermore, to mitigate the issues of over-smoothing and distortion and obtain finer reconstruction results step-by-step, we design a multi-stage training strategy, as illustrated by Algorithm 1. Specifically, we divide the training process into four stages, i.e., $stage=stage_{i},when{{\color[rgb]{1,1,1}1}}epoch\in\left[E_{i-1},E_{i}\right]|_{i=1}^{i=4}$. In summary, our IGIT-Net reconstructs the precise k-space spectrum and MRI images progressively by optimizing different loss $L_{stage_{j}}$ at various stages as follows:
$L_{stage_{j}}={\textstyle\sum_{i=1}^{i=j}L_{i}}$ (20)
4 Experiments
4.1 Implementation Details
4.1.1 Datasets
We use two real single-coil k-space datasets, CC359 (Warfield et al., 2004) and fastMRI (Zbontar et al., 2018), one simulated k-space dataset, IXI (Orhaug and Forssell, 2021), and one real multi-coil k-space dataset, SKM-TEA (Desai et al., 2022), to validate the effectiveness of our method.
@@ 4.1.1 Datasets
CC359 dataset The brain MR raw dataset—the Calgary-Campinas dataset, which comes from a clinical MR scanner (Discovery MR750; GE Healthcare, Waukesha, WI, USA)—is used for validation. $4129$ slices from $25$ volumes are selected to form the training set, and the validation set is composed of $1650$ slices from $10$ other volumes.
@@
fastMRI dataset The fastMRI dataset is currently the largest open-source MRI dataset with raw fully sampled k-space data. The single-coil knee dataset which consists of $973$ training volumes ($98772$D slices) and $199$ validation volumes ($6140$ 2D slices) is employed in the experiments.
@@
IXI dataset The IXI dataset consists of $574$ brain MRI volumes collected from three different hospitals in London. After pre-processing, $46226$ 2D slices from $368$ volumes are used for training, $11562$ 2D slices from $92$ volumes are used for validation, and $14315$ 2D slices from $114$ volumes are used for testing. Since IXI does not contain original k-space data, we obtain the simulated k-space via FFT.
SKM-TEA dataset The SKM-TEA raw data provides 155 multi-coil T2-weighted knee MRI scans, and each provides approximately 160 cross-sectional knee images with the matrix of size $512\times 512$. Following the well-established ReconFormer  (Guo et al., 2023), we used 124, 10, and 21 volumes for training, validation, and testing, respectively.
4.1.2 K-space undersampling
@@
In our experiments, two different types of masks are tested, i.e., 1D Cartesian mask and 2D Gaussian mask. For each mask, $20\%$ and $40\%$ undersampling rates are tested. Following (Zbontar et al., 2018), all undersampling masks are generated by first including some number of adjacent lowest frequency k-space lines or points to provide a fully-sampled k-space region. Specifically, for the 1D Cartesian undersampling mask, the fully-sampled central region includes $8\%$ of all k-space lines; for the 2D Gaussian undersampling mask, $16\%$ of all k-space points are included. Then, the remaining k-space lines (for 1D Cartesian mask) or points (for 2D Gaussian mask) are included at random, with the probability set so that, on average, the undersampling mask achieves the desired acceleration factor.
4.1.3 Baselines and Training Details
@@
We implement our model in PyTorch using Adam with an initial learning rate of $31e\-/$ 3. $(E_{0},E_{1},E_{2},E_{3},E_{4})$ for multi-stage training are set as $(0,20,60,100,200)$.
We compare our model with the following algorithms. Zero-Filling (ZF); ResNet(He et al., 2016); UNet(Zbontar et al., 2018); RefineGAN(Quan and Jeong, 2018), a generative adversarial model; SwinMR(Huang et al., 2022), a swin Transformer based method; SwinGAN(Zhao et al., 2023), which combines the swin transformer and generative adversarial network for fast MRI; ReconFormer(Guo et al., 2023), which designs a locally pyramidal but globally columnar transformer for multi-scale modeling; KIKI-Net(Eo et al., 2018), which combines 4 CNNs to operate on k-space and image sequentially; D5C5(Schlemper et al., 2018), a deep cascaded network for fast MRI. For fair comparison, we retrain these models using their default parameter settings.
@@
Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Normalized Mean Squared Error (NMSE), and Learned Perceptual Image Patch Similarity (LPIPS) are utilized for quantitative evaluation.
4.2 Experimental Results
4.2.1 Single-coil datasets
@@
The quantitative results of comparisons on single-coil datasets including CC359, fastMRI, and IXI datasets are shown in Table 1, Table 2, and Table 3, respectively. The best result in each column is highlighted in bold. Based on these results, we can find that: (1) Accelerated sampling leads to zero-filled MRI images with extremely poor quality; (2) Classical deep learning methods, such as UNet and ResNet, reduce image artifacts to some extent, but the restoration results remain suboptimal; (3) Current state-of-the-art methods, such as RefineGAN, SwinMR, and ReconFormer, have significantly improved image reconstruction quality, but their performance degrades on challenging tasks with high acceleration rates (sampling rate = 20%). (4) Our method consistently achieves the best performance with significant improvements on all datasets under different sampling scenarios. For example, as shown in Table 1, our method shows the superiority of $1.08$ dB and $1.69$ dB in PSNR over ReconFormer under 20% and 40% 2D Gaussian mas sampling masks on the CC359 dataset, respectively. As shown in Table 3, our model outperforms ReconFormer by $0.51$ dB and $0.29$ dB in PSNR under 20% and 40% 1D Cartesia masks on the IXI dataset, respectively.
4.2.2 Multi-coil dataset
We further validate the effectiveness of our method on the multi-coil dataset SKM-TEA. The results in Table 4 show that our method still exhibits significant superiority over other competitors on the real multi-coil k-space dataset.
4.2.3 Experiments on masks
This experimental study aims to evaluate the robustness of our IGKR-Net using different undersampling trajectories. Five 1D Cartesian undersampling trajectories including 1D Cartesian $10\%$, 1D Cartesian $20\%$, 1D Cartesian $30\%$, 1D Cartesian $40\%$ , and 1D Cartesian $50\%$, as well as five 2D Random undersampling trajectories including 2D Random $10\%$, 2D Random $20\%$, 2D Random $30\%$, 2D Random $40\%$, and 2D Random $50\%$ are applied in this experiment. This experiment is conducted using the CC359 dataset.
The quantitative results are shown in Table 5. According to the results, we can fine that: (1) When using the same type of mask, e.g., 1D Cartesian mask, our IGKR-Net consistently demonstrates significant improvements across different sampling rates compared to other methods. For example, our model outperforms refineGAN by $3.66$ dB and $0.21$ dB under 1D Cartesian mask at $50\%$ and $40\%$ sampling rates, respectively. (2) As shown in Figure.6, our method’s performance under 2D masks (denoted as solid lines) even consistently surpasses that of other methods under 1D masks (denoted as dashed lines ) across different sampling rates, despiting 1D mask being recognized as a simpler task. (3) RefineGAN achieves satisfactory results at higher sampling rates, while its performance significantly declines under more challenging lower rates. Compared to RefineGAN, our method shows robustness across various masks.
@@ 4.2.3 Experiments on masks
Figure 6: Results of the mask experiments. The X-axis denotes the sampling ratio, while the Y-axis corresponds to various metrics.
Table 5: Performance comparison of MRI reconstruction under different undersampling masks on the multi-coil SKM-TEA.
@@
ratio Method PSNR↑ SSIM↑ NMSE↓ PSNR↑ SSIM↑ NMSE↓
Unet 25.72 0.8010 0.0352 26.50 0.7909 0.0301
RefineGAN 20.80 0.5677 0.3890 26.27 0.7363 0.0327
10% Ours 27.11 0.8262 0.0267 28.96 0.8424 0.0181
Unet 29.34 0.8756 0.0156 28.34 0.8417 0.0197
RefineGAN 29.96 0.8888 0.0303 30.10 0.8554 0.0324
20% Ours 33.06 0.9234 0.0069 34.26 0.9258 0.0058
UNet 30.15 0.8940 0.0127 29.73 0.8738 0.0147
RefineGAN 33.84 0.9286 0.0152 33.62 0.9110 0.0162
30% Ours 35.35 0.9530 0.0039 38.38 0.9600 0.0024
UNet 31.73 0.9189 0.0090 30.59 0.8935 0.0122
RefineGAN 38.57 0.9656 0.0058 36.58 0.9396 0.0092
40% Ours 38.78 0.9703 0.0020 41.22 0.9753 0.0013
Unet 34.08 0.9442 0.0053 32.06 0.9160 0.0087
RefineGAN 39.68 0.9746 0.0047 38.84 0.9571 0.0056
50% Ours 43.34 0.9844 0.0008 46.04 0.9875 0.0005
4.2.4 Visualization Results
The qualitative comparison results are shown in Figure.5, including the reconstructed MRI images and absolute error maps. As observed, reconstructed MRI image using zero-filling (ZF) at high acceleration rates leads to significant aliasing artifacts and loss of anatomical details. In comparison to ZF, deep learning methods based on CNNs such as ResNet and UNet show improvement in reconstruction but still suffer from severe edge blurring and substantial detail loss. SwinGAN and ReconFormer partially alleviate these issues, yet in challenging tasks with high acceleration rates, these methods lose some crucial anatomical details. In contrast, our IGKR-Net demonstrates robustness to sampling patterns and acceleration rates. Leveraging information from the image domain to guide the recovery in k-space, our method better preserves important anatomical details, as indicated by the enlarged boxes and yellow ellipses.
@@ 4.2.4 Visualization Results
As shown in Figure.7, as the k-space sampling ratio increases, the performance of different methods steadily improves. It is noteworthy that in challenging tasks with low sampling rates, e.g., 1D or 2D $10\%$, other methods obtain extremely poor performance, while our method can still reconstruct relatively clear contour edges. Overall, our IGKR-Net achieved a higher reconstruction quality compared to other methods under various undersampling trajectories.
Table 6: Quantitative results of the ablation studies using 1D Cartesian mask (sampling ratio = 20%) on the CC359 dataset.
Model LRIT HRIT IDGM TARM PSNR↑ SSIM↑ NMSE↓ LPIPS↓
(a) $\checkmark$ $\checkmark$ $\checkmark$ 31.65 0.8996 0.0098 0.0552
(b) $\checkmark$ $\checkmark$ $\checkmark$ 28.56 0.7362 0.0535 0.1027
(c) $\checkmark$ $\checkmark$ $\checkmark$ 30.42 0.8826 0.0120 0.0612
(d) $\checkmark$ $\checkmark$ $\checkmark$ 31.97 0.9091 0.0086 0.0440
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 33.06 0.9234 0.0069 0.0362
Figure 7: Visual comparison results of the mask experiments.
4.3 Ablation Studies and Further Analysis
4.3.1 Efficacy of Key Components
To validate the effectiveness of the Key Components within our proposed method, we conduct a breakdown ablation experiment, the results are shown in Table 6. We can observe that: (a) When removing the low- to high-resolution progressive learning strategy, i.e., querying the unsampled k-values directly using high-resolution coordinates, the model produces a significant performance degradation from PSNR $33.06$ to $31.65$. Direct recovery of high-resolution k-space from overly undersampled data leads to oversmoothing, which results in poorer reconstruction results. This illustrates the effectiveness of our low- to high-resolution multi-stage learning strategy. (b) Disabling the HRIT module, i.e., directly up-sampling the low-resolution outputs obtained by LRIT as the reconstruction results, resulted in a severe decrease of (PSNR, SSIM) by $(4.50,0.1872)$, respectively. This suggests that using the dense k-space obtained from LRIT to further construct a more accurate latent space, and querying the k-space by high-resolution coordinates is important for effective MRI reconstruction. (c) When removing IDGM from the network and using the k-space spectrum outputted from the LRIT for HRIT, the performance decreased significantly, which demonstrates that our proposed IDGM can mine information from undersampled MRI images to obtain more informative outputs. (d) Disabling the TARM led to a $1.09$ drop in PSNR and $0.0143$ decrease in SSIM, demonstrating its importance in refining the reconstruction outputs in image domain, thereby achieving accurate results.
4.3.2 Efficacy of Recovering K-space